OpenStudy (anonymous):

determine weather the graphs of the pair of lines are parallel or undefined x+10=y y-x=-6 so x+10=y is? or is it undefined so y-x=-6 is? or is it undefined

OpenStudy (anonymous):

y=x+10 y=x-6 the coefficient in a front of x are equal 1, then lines have the same slope, so - parallel